Sketch the graph of f(x) given the following: lim f(x) = 2 f(-1) =1 f(1) =1 lim f(x) = DNE 1--1 ->
A function f is continuous at a number a if lim y(x) = f(a) ca requires the following 3 things:
Answer #1

lim f(x) = 2 f(-1)= 1, f(1) =1, lim 50 = DNE ) x-1 -> 1 2 2x -27 lim f(1) = -2 *رود lim f(31) = 1 ) i lim f(x) = DINE 3-1A function at q ol Point is Continuous ib a at The number lim f(a)= f(a) - G Requires following 3 things ① The function shoul